OpenStudy (anonymous):

What's the exact value of cos(pi/12) using the half-angle formula?

zepdrix (zepdrix):

Hey! \(\Large\bf \color{#DD4747 }{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

\[\Large\bf\sf \cos\left(\frac{x}{2}\right)\quad=\quad \pm\sqrt{\frac{1+\cos x}{2}}\]Since pi/12 is in the first quadrant, we only want the `positive` square root.

zepdrix (zepdrix):

\[\Large\bf\sf \cos\left(\frac{\pi}{12}\right)\quad=\quad \cos\left(\frac{\pi}{6\cdot 2}\right)\quad=\quad \cos\left(\frac{\pi/6}{2}\right)\] \[\Large\bf\sf =\quad \sqrt{\frac{1+\cos \pi/6}{2}}\] Lemme know if that second to last step was confusing. I did a little trick with fractions.